A 45.3 - cm radius wheel is uniformly speeding up from 146 rev/min to 621 rev/min within 26.5 s. Find the tangential acceleration.


1
Expert's answer
2021-03-19T06:48:00-0400

Here revolution per min is given

"N_1=" 146 rev/min , "\omega_1= \frac{2\pi N_1}{60}= \frac{2\times 3.14\times 146}{60}=15.28 \frac{rad}{s}"


"N_2= \omega_2= \frac{2\pi N_2}{60}= \frac{2\times 3.14\times 621}{60}=64.998 \frac{rad}{s}"


t=26.5 s,

using first equation of rotation motion

"\omega_2=\omega_1+ \alpha t"


"\alpha=1.876 \frac{rad}{s^2}"

"\alpha" = angular acceleration


tangential acceleration,

"a= r \alpha=0.85 \frac{m}{s^2}"